Use the simple interest formula to determine the missing value.
p=​$964.55​, r=7.5​%, t=​?, i=​$434.05

Answer by fractalier(6550)

Since I = Prt, we can find t by dividing by Pr, so that
t = I/Pr = 434.05/(964.55*.075) = 6 years